# OODFace: Benchmarking Robustness of Face Recognition under Common Corruptions and Appearance Variations

### v1

@@ Abstract
With the rise of deep learning, facial recognition technology has seen extensive research and rapid development. Although facial recognition is considered a mature technology, we find that existing open-source models and commercial algorithms lack robustness in certain real-world Out-of-Distribution (OOD) scenarios, raising concerns about the reliability of these systems. In this paper, we introduce OODFace, which explores the OOD challenges faced by facial recognition models from two perspectives: common corruptions and appearance variations. We systematically design 30 OOD scenarios across 9 major categories tailored for facial recognition. By simulating these challenges on public datasets, we establish three robustness benchmarks: LFW-C/V, CFP-FP-C/V, and YTF-C/V. We then conduct extensive experiments on 19 different facial recognition models and 3 commercial APIs, along with extended experiments on face masks, Vision-Language Models (VLMs), and defense strategies to assess their robustness. Based on the results, we draw several key insights, highlighting the vulnerability of facial recognition systems to OOD data and suggesting possible solutions. Additionally, we offer a unified toolkit that includes all corruption and variation types, easily extendable to other datasets. We hope that our benchmarks and findings can provide guidance for future improvements in facial recognition model robustness.
@@ 1 Introduction
(Problem statement) Since FR has become a well-established technology, are FR models capable of handling all complex real-world scenarios?
@@
In real-life scenarios, users still encounter issues, such as failing to unlock apartment access during snowy weather or unsuccessful facial recognition for apps that have not been used for a while [2]. To explore the causes, we conduct a simple experiment by adding light Gaussian noise to facial images (potentially introduced during data processing or transmission), as shown in Fig. 1. Although humans can easily recognize the person in both images, we find that state-of-the-art open-source models [9] and the commercial API of Tencent experience a significant drop in performance, with the detection success rate dropping to 82.2%. The API fails to detect faces in 48.9% of the images, disrupting the normal functioning of the FR system.
@@
In this paper, we explore the OOD challenges faced by FR models from the perspectives of common corruptions and appearance variations. We systematically design 30 common OOD scenarios tailored for FR to rigorously assess the OOD robustness of current models. Common corruptions are classified into five categories: Lighting & Weather, Sensor Failures, Motion Errors, Data Processing Corruptions, and Object Occlusion, covering 20 subclasses. Appearance variations are grouped into four categories: Age, Facial Expression, Head Pose, and Accessories, comprising 10 subclasses, covering most real-world disturbances (see Fig. 2). Many of these scenarios, such as Facial Distortion, Random Occlusion, Age, and Accessories, are designed specifically for FR and have not been previously explored by [52, 28, 2, 56, 36]. Following [20], each scenario includes five severity levels, resulting in 150 unique corruptions and variations. By applying these scenarios to standard FR datasets—LFW [22], CFP-FP [42], and YTF [49]—we establish three comprehensive benchmarks for robustness evaluation: LFW-C/V, CFP-C/V, and YTF-C/V, representing common corruptions and appearance variations, respectively. We hope these large-scale OOD datasets can serve as standard benchmarks for fair and comprehensive evaluation of OOD robustness in FR models, facilitating future research in the field.
We conduct extensive experiments to compare the OOD robustness of existing FR models. Specifically, we evaluate 19 open-source models and 3 commercial APIs, covering a variety of loss functions, backbone networks, and model sizes, across the LFW-C/V, CFP-FP-C/V, and YTF-C/V benchmarks. Based on the evaluation results, we find that: 1) Common corruptions robustness is uncorrelated with clean data performance, whereas appearance variations robustness shows the opposite trend; 2) FR models show significant performance drops under common corruptions, with the largest impact from Data & Processing; 3) different FR models show varying sensitivity to different types of OOD scenarios. More discussions are provided in Sec. 5. Additionally, we explore the effectiveness of face masks under OOD conditions and investigate leveraging the generalization capabilities of Vision-Language Models to address OOD challenges. Finally, we examine 10 robust models using input transformation [50, 51] and adversarial training [32, 53] as potential defense for improving robustness in Appendix F. However, the observed robustness gains are limited. Therefore, despite their widespread use in critical applications, FR models still exhibit vulnerabilities, and enhancing their robustness remains an open challenge.
@@ 2.1 Face Recognition
Facial recognition, a key task in computer vision, has made significant progress in recent years. Early models like DeepFace [44] and FaceNet [41] were the first to demonstrate the power of deep learning for FR tasks. Later, models such as SphereFace [30], CosFace [47], and ArcFace [10], which leverage angular margin-based loss functions, further enhanced feature discriminability. AdaFace [23] introduced an adaptive margin function to adjust sample weighting based on image quality, improving performance on low-quality datasets. TransFace [8] focuses on FR by employing a patch-level data augmentation strategy (DPAP) and an entropy-based hard sample mining (EHSM), which preserves facial structure while enhancing sample diversity. The latest TopoFR [9] model incorporates topology alignment and hard sample mining, effectively preserving data structure and improving generalization.
@@ 3 OOD in Face Recognition
We introduce the categories of common corruptions and appearance variations in Sec. 3.1 and Sec. 3.2, respectively, and provide implementation details in Appendix. A.
@@ 3.2 Appearance Variations
Accessory & Makeup: In everyday life, people often wear various accessories that can obscure or alter facial features [36]. Based on [26], we add accessories to the images to simulate real-world accessories occlusions Makeup, especially common in social settings, can visually alter facial features by changing eyebrow shapes, eyeliner, blush, and lip color. We follow [25] to add diverse makeup styles.
Variation Levels: Similar to common corruptions, we define five severity levels for appearance variations (see Fig. 4) For age adjustments, the levels indicate the degree of age reduction or increase. For facial expressions, the levels represent the extent of mouth and eye movements. Head pose levels denote the degree of facial rotation. For accessory additions, the levels are defined as follows: glasses (clear lenses), bang occlusion, glasses (reflective lenses), glasses (clear lenses) + bang, and glasses (reflective lenses) + bang. For the makeup variation, we adopt the styles defined in [25]: Japanese style, Korean style, Retro style, Flashy style, and Smoky-eyes style. Each level represents an increase in perceived appearance interference.
@@ 4 OOD Benchmarks
To thoroughly evaluate the robustness of FR models, we establish multiple robustness benchmarks using widely adopted datasets: LFW [22], CFP-FP [42], and YTF [49]. For each dataset, we create two benchmark versions: one for common corruptions (denoted as LFW-C, CFP-C, YTF-C) and one for appearance variations (denoted as LFW-V, CFP-V, YTF-V). Additionally, we offer a unified toolkit that includes all corruptions and variations, easily extendable to other datasets. Below, we detail the datasets, evaluation metrics, and FR models used in our benchmarks.
@@ 4.1 LFW-C/V
We first conduct experiments using the LFW [22] dataset, a widely used benchmark in FR. LFW consists of 5,749 identities and 13,233 images, forming 6,000 pairs of face images. We preprocess the entire LFW dataset using MTCNN [54], obtaining aligned images. We then apply 20 types of common corruptions and 10 types of appearance variations. Each includes five levels of severity [20].
Common Corruption Dataset LFW-C. We design LFW-C, a version of the LFW dataset that incorporates common corruptions, to comprehensively evaluate the robustness of FR models under various types and severities of corruption. For each model, we first obtain its performance on the original LFW dataset, denoted as $\mathrm{Acc}_{clean}$. We then re-evaluate the model’s performance under each corruption type $c$ and severity level $s$ in LFW-C, denoted as $\mathrm{Acc}_{c,s}$.
@@
Appearance Variations Dataset LFW-V. We also design LFW-V, an appearance editing version of the LFW. Similarly, for the LFW-V, we first obtain the $\mathrm{Acc}_{clean}$ and then evaluate the model’s performance under each variation category $v$ and severity level $s$, denoted as $\mathrm{Acc}_{v,s}$.
@@ 4.2 CFP-C/V
The CFP (Celebrities in Frontal-Profile) [42] dataset consists of frontal and profile images of celebrities. It includes clear frontal and side-view face images, with a total of 500 identities and 7,000 pairs. Following a similar procedure as LFW, we first align all images using MTCNN [54], then apply designed 20 types of common corruptions and 10 types of appearance variations.
Common Corruption Dataset CFP-C. Similarly, CFP-C introduces 20 types of common corruptions into the CFP validation set. These corruptions are applied with five different levels of severity. For each model, we also first compute $\mathrm{Acc}_{clean}$, and then evaluate the model under each corruption type $c$ and severity level $s$, denoted as $\mathrm{Acc}_{c,s}$. We follow the Equation 1 to calculate $\mathrm{Acc_{cor}}$ and Equation 2 to compute the RCE in a similar manner.
Appearance Variations Dataset CFP-V. We also design CFP-V, an appearance variations version based on the CFP dataset, and calculate $\mathrm{Acc}_{clean}$, $\mathrm{Acc}_{v,s}$, $\mathrm{Acc_{var}}$, RVE in a comparable format with LFW-V.
@@ 5 Benchmarking Results and Insights
We present the evaluation results on LFW-C in Sec. 5.1, and LFW-V in Sec. 5.2, and leave the results on CFP-C/V and YTF-C/V in Appendix C and Appendix D. In Sec. 5.3, we discuss commercial API evaluation, followed by extended experiments on face masks in Sec. 5.4 and VLMs in Sec. 5.5. Key findings are summarized by insight block.
@@ 5.1 Common Corruptions Evaluation Results
Tab. 2 shows the robustness evaluation of 19 FR models on LFW-C, categorized into Open-source Model Eval, Architecture Eval, and Loss Function Eval. We report the average performance across five levels for each category. It is evident that natural robustness does not strongly correlate with $\mathrm{Acc}_{clean}$. For example, models with high $\mathrm{Acc}_{clean}$ (e.g., TopoFR [9]) does not achieve the high $\mathrm{Acc_{cor}}$. In Fig. 5, we further illustrate the Relative Corruption Error (RCE) for each model across different corruption categories. Based on these evaluations, we provide the following analysis:
@@
Insight 1: FR models experience significant performance drops under common corruptions, with the largest impact from Data & Processing, such as Gaussian noise and salt-and-pepper noise.
Comparison of FR Models. While all models experience performance decline, different models show varying sensitivities to specific corruptions. For instance, FaceNet [7] demonstrates robustness against Sensor and Motion corruptions, possibly due to its use of Triplet Loss. AdaFace [23] achieves the best average FR accuracy across all corruptions, attributed to its adaptive margin loss function, which adjusts sample weighting based on image quality, approximated by the feature norm. Meanwhile, ArcFace [23] achieves the highest accuracy in Data & Processing, benefiting from its Angular Margin Loss, which provides robust features even under low-resolution or noisy sensor data.
Model Architecture Comparison. MobileFace [7] achieves the best average robustness with the fewest parameters (1.2M), particularly excelling in the Data & Processing and Occlusion. ResNet50 [19] ranks second in average robustness, showing strong performance in the Lighting & Weather, and Sensor categories.
Loss Function Comparison. Among the loss functions, LMCL [47] achieves the best average results, especially in the Lighting & Weather. This can be attributed to its optimization of inter-class boundaries in the feature space, enhancing model discrimination for similar samples. Additionally, AM-Softmax [46] shows superior robustness in the Data & Processing category.
Insight 2: Different models show varying sensitivity to different types of OOD scenarios. Some models are more robust to specific OOD types but vulnerable to others, indicating that model architecture, loss functions, and training procedures significantly influence performance under specific conditions.
Figure 5: RCE results on LFW-C.
@@ 5.2 Appearance Variations Evaluation Results
Tab. 3 presents the variation robustness evaluation of 19 FR models on LFW-V. Overall, the impact of appearance variations is smaller than that of corruptions. Unlike corruptions, robustness against appearance variations is largely correlated with the model’s clean accuracy, consistent with observations in [12, 20]. In Fig. 6, we further illustrate the Relative Variation Error (RVE) for each model across variations categories. We provide the following analysis:
@@
Comparison of Variations Types. As shown in Tab. 3 and Fig. 6, appearance variations also degrade model performance, but their impact is generally less severe than corruptions, with the highest RVE around 4%. This could be because, unlike corrupted data, FR models are trained with data that includes redundancy for changes such as age, accessories, and others, making these variations closer to in-distribution data. Additionally, the Accessories have the least overall impact, likely because appearance changes like makeup are superficial and do not alter key facial features. In contrast, age variations have the most significant effect, as aging causes notable changes in facial features, such as increased wrinkles and skin laxity.
Insight 3: The impact of appearance variations is smaller, likely due to redundancy in training data that improves robustness to these changes.
@@
Loss Function Comparison. CosFace-IR and ArcFace-IR are tied as the top performers, both achieving an average accuracy of 97.17%. This can be attributed to the design of LMCL [47] and ArcFace [10] losses, which enhance robustness to complex variations. Additionally, their high clean accuracy suggests that these models effectively learn facial representations, enabling them to maintain capabilities when facing appearance changes.
Insight 4: Model robustness against common corruptions does not fully correlate with clean data performance. Even models that excel on clean data may struggle with certain corruptions. In contrast, robustness to appearance variations is strongly correlated with clean data performance.

### v2

@@ Abstract
With the rise of deep learning, facial recognition technology has seen extensive research and rapid development. Although facial recognition is considered a mature technology, we find that existing open-source models and commercial algorithms lack robustness in certain complex Out-of-Distribution (OOD) scenarios, raising concerns about the reliability of these systems. In this paper, we introduce OODFace, which explores the OOD challenges faced by facial recognition models from two perspectives: common corruptions and appearance variations. We systematically design 30 OOD scenarios across 9 major categories tailored for facial recognition. By simulating these challenges on public datasets, we establish three robustness benchmarks: LFW-C/V, CFP-FP-C/V, and YTF-C/V. We then conduct extensive experiments on 19 facial recognition models and 3 commercial APIs, along with extended physical experiments on face masks to assess their robustness. Next, we explore potential solutions from two perspectives: defense strategies and Vision-Language Models (VLMs). Based on the results, we draw several key insights, highlighting the vulnerability of facial recognition systems to OOD data and suggesting possible solutions. Additionally, we offer a unified toolkit that includes all corruption and variation types, easily extendable to other datasets. We hope that our benchmarks and findings can provide guidance for future improvements in facial recognition model robustness.
@@ 1 Introduction
(Problem statement) Since FR has become a well-established technology, are FR models capable of handling all complex scenarios?
@@
In real-life situations, users may encounter issues, such as failing to unlock apartment access during snowy weather or unsuccessful facial recognition for apps that have not been used for a while [3]. To explore the causes, we conduct a simple experiment by adding light Gaussian noise to facial images (potentially introduced during data processing or transmission), as shown in Fig. 1. Although humans can easily recognize the person in both images, we find that state-of-the-art open-source models [10] and the commercial API of Tencent experience a significant drop in performance, with the detection success rate dropping to 82.2%. The API fails to detect faces in 48.9% of the images, disrupting the normal functioning of the FR system.
@@
In this paper, we explore the OOD challenges faced by FR models from the perspectives of common corruptions and appearance variations. We systematically design 30 common OOD scenarios tailored for FR to rigorously assess the OOD robustness of current models. Common corruptions are classified into five categories: Lighting & Weather, Sensor Failures, Motion Errors, Data Processing Corruptions, and Object Occlusion, covering 20 subclasses. Appearance variations are grouped into four categories: Age, Facial Expression, Head Pose, and Accessories, comprising 10 subclasses, covering most real-world disturbances (see Fig. 2). Following [22], each scenario includes five severity levels, resulting in 150 unique corruptions and variations. Many of these scenarios, such as Facial Distortion, Random Occlusion, Age, and Accessories, are designed specifically for FR and have not been previously explored by [56, 30, 3, 60, 39], which also investigate OOD challenges but focus on specific OOD types with a limited evaluation scope and fewer FR models (Detailed comparison in Appendix. A.3, Tab. A.3). By applying these scenarios to standard FR datasets—LFW [24], CFP-FP [45], and YTF [53]—we establish three comprehensive benchmarks for robustness evaluation: LFW-C/V, CFP-C/V, and YTF-C/V, representing common corruptions and appearance variations, respectively. We hope these large-scale OOD datasets can serve as standard benchmarks for fair and comprehensive evaluation of OOD robustness in FR models, facilitating future research in the field.
We conduct extensive experiments to compare the OOD robustness of existing FR models. Specifically, we evaluate 19 open-source models and 3 commercial APIs, covering a variety of loss functions, backbones, and model sizes, across the LFW-C/V, CFP-FP-C/V, and YTF-C/V benchmarks. Based on the evaluation results, we find that: 1) Common corruptions robustness is uncorrelated with clean data performance, whereas appearance variations robustness shows the opposite trend; 2) FR models show significant performance drops under common corruptions, with the largest impact from Data & Processing; 3) different FR models show varying sensitivity to different types of OOD scenarios. More discussions are provided in Sec. 5. Additionally, we explore the effectiveness of physical face masks under OOD conditions. Finally, we examine 10 robust models using input transformation [54, 55], adversarial training [35, 57] and 3 restoration methods [51, 63, 31] based on GANs, Transformers, and Diffusion models as potential defense, and investigate leveraging the generalization capabilities of Vision-Language Models to address OOD challenges. However, the observed robustness gains are limited. Therefore, despite their widespread use in critical applications, FR models still exhibit vulnerabilities, and enhancing their robustness remains an open challenge.
@@ 2.1 Face Recognition
Face recognition, a key task in computer vision, has made significant progress in recent years. Early models like DeepFace [47] and FaceNet [44] are the first to demonstrate the power of deep learning for FR tasks. Later, models such as SphereFace [33], CosFace [50], and ArcFace [11], which leverage angular margin-based loss functions, further enhanced feature discriminability. AdaFace [25] introduced an adaptive margin function to adjust sample weighting based on image quality, improving performance on low-quality datasets. TransFace [9] focuses on FR by employing a patch-level augmentation strategy (DPAP) and an entropy-based hard sample mining (EHSM), which preserves facial structure while enhancing sample diversity. The latest TopoFR [10] model incorporates topology alignment and hard sample mining, effectively preserving data structure and improving generalization.
@@ 3 OOD in Face Recognition
We introduce the categories of common corruptions and appearance variations in Sec. 3.1 and Sec. 3.2, respectively, with more details provided in Appendix. A. These include implementation details for corruptions and variations (A.1 & A.2), a comparison with related work (A.3), and a discussion on the naturalness of OOD synthesis (A.4).
@@ 3.2 Appearance Variations
Accessory & Makeup: In everyday life, people often wear various accessories that can obscure or alter facial features [39]. Based on [28], we add accessories to the images to simulate real-world accessories occlusions. Makeup, especially common in social settings, can visually alter facial features by changing eyebrow shapes, eyeliner, blush, and lip color. We follow [27] to add diverse makeup styles. Variation Levels. Similar to common corruptions, we define five severity levels for appearance variations (see Fig. 4) , with detailed level settings provided in Appendix A.3.
Naturalness of OOD Synthesis. Our synthesis method can be regarded as closely approximating real-world effects, with evaluation and analysis detailed in Appendix A.4.
@@ 4 OOD Benchmarks
To thoroughly evaluate the robustness of FR models, we establish robustness benchmarks using widely adopted datasets: LFW [24], CFP-FP [45], and YTF [53]. For each dataset, we create two benchmark versions: one for common corruptions (denoted as -C) and one for appearance variations (denoted as -V). Below, we detail the datasets, evaluation metrics, and FR models used in our benchmarks.
@@ 4.1 LFW-C/V
We first conduct experiments using the LFW [24] dataset, a widely used benchmark in FR. LFW consists of 5,749 identities and 13,233 images, forming 6,000 pairs of face images. We preprocess the entire LFW dataset using MTCNN [58], obtaining aligned images. We then apply 20 types of common corruptions and 10 types of appearance variations. Each includes five levels of severity.
Corruption Dataset LFW-C. To comprehensively evaluate the robustness of FR models under various types of corruption. For each model, we first obtain its performance on the original LFW dataset, denoted as $\mathrm{Acc}_{clean}$. We then re-evaluate the model’s performance under each corruption type $c$ and severity level $s$ in LFW-C, denoted as $\mathrm{Acc}_{c,s}$.
@@
Variations Dataset LFW-V. Similarly, we obtain the $\mathrm{Acc}_{clean}$ and evaluate the model’s performance under each variation category $v$ and severity level $s$, denoted as $\mathrm{Acc}_{v,s}$.
@@ 4.2 CFP-C/V
The CFP [45] dataset consists of frontal and profile images of celebrities. It includes clear frontal and side-view face images, with a total of 500 identities and 7,000 pairs. Following a similar procedure as LFW, we first align all images using MTCNN [58], then apply designed 20 types of common corruptions and 10 types of appearance variations.
Corruption Dataset CFP-C. Similarly, CFP-C introduces 20 types of corruptions into the CFP validation set. These corruptions are applied with five levels of severity. For each model, we also first compute $\mathrm{Acc}_{clean}$, and evaluate it under each corruption type $c$ and severity level $s$, denoted as $\mathrm{Acc}_{c,s}$. We follow the Equation 1 to calculate $\mathrm{Acc_{cor}}$ and Equation 2 to compute the RCE in a similar manner.
Variations Dataset CFP-V. We also design CFP-V, a variations version based on the CFP, and calculate $\mathrm{Acc}_{clean}$, $\mathrm{Acc}_{v,s}$, $\mathrm{Acc_{var}}$, RVE in a comparable format with LFW-V.
@@ 5 Benchmarking Results and Insights
We present the evaluation results on LFW-C in Sec. 5.1, and LFW-V in Sec. 5.2, and leave the results on CFP-C/V and YTF-C/V in Appendix C and Appendix D. In Sec. 5.3, we discuss commercial API evaluation, followed by extended physical experiments on face masks in Sec. 5.4.
@@ 5.1 Common Corruptions Evaluation Results
Tab. 2 shows the robustness evaluation of 19 FR models on LFW-C, categorized into Open-source Model Eval, Architecture Eval, and Loss Function Eval. We report the average performance across five levels for each category. It is evident that model robustness does not strongly correlate with $\mathrm{Acc}_{clean}$. For example, models with high $\mathrm{Acc}_{clean}$ (e.g., TopoFR [10]) does not achieve the high $\mathrm{Acc_{cor}}$. In Fig. 5, we further illustrate the Relative Corruption Error (RCE) for each model across corruption categories. Based on these evaluations, we provide the following analysis:
Figure 5: RCE results on LFW-C.
@@
Insight 1: FR models suffer significant performance degradation under corruptions, with Data & Processing causing the most severe impact. This occurs as noise disrupts the feature space distribution, shifting the decision boundary of linear classifiers, and the destructiveness of noise corruption also stems from the model’s over-reliance on high-frequency phase, which is linked to the frequency response characteristics of FR architectures, leading to degraded cross-layer feature fusion. (Virtualized in Appendix Fig. A.1)
Comparison of FR Models. While all models experience performance decline, different models show varying sensitivities to specific corruptions. For instance, FaceNet [8] demonstrates robustness against Sensor and Motion corruptions. ArcFace [25] achieves the highest accuracy in Data & Processing, benefiting from its Angular Margin Loss, which provides robust features even under low-resolution or noisy sensor data. Meanwhile, AdaFace [25] achieves the best average FR accuracy across all corruptions, attributed to its adaptive margin loss function.
Comparison of Architecture and Loss Function. MobileFace [8] achieves the best average robustness with the fewest parameters (1.2M), particularly excelling in the Data & Processing and Occlusion. ResNet50 [21] ranks second in average robustness, showing strong performance in the Lighting & Weather, and Sensor categories. Among the loss functions, LMCL [50] achieves the best average results. This can be attributed to its optimization of inter-class boundaries in the feature space, enhancing discrimination for similar samples. Additionally, AM-Softmax [49] shows superior robustness in the Data & Processing category.
Insight 2: Different models exhibit varying sensitivity to OOD scenarios, with some being robust to specific OOD types while vulnerable to others. This highlights the critical role of architecture, loss functions, and training procedures in shaping model performance under different conditions. Loss functions induce distinct decision boundary geometries. For instance, AdaFace leverages spherical compression with adaptive feature norm adjustment, mitigating the local perturbation sensitivity observed in losses like Triplet Loss. Mainstream datasets are curated with limited diversity, lacking spatio-temporal noise modeling in real-world conditions. The feature representation capacity of models is constrained by data quality.
@@ 5.2 Appearance Variations Evaluation Results
Comparison of Variations Types. As shown in Tab. 3 and Fig. 6, appearance variations also degrade model performance, but their impact is generally less severe than corruptions, with the highest RVE around 4%. Unlike corrupted data, FR models are trained with data that includes redundancy for changes such as age, accessories, and others. Additionally, Accessories exhibit the least overall degradation, likely because variations such as makeup are superficial and do not significantly alter key facial structures. In contrast, Age variations induce the most substantial impact, as aging introduces prominent facial changes, including increased wrinkles and skin laxity.
@@
Insight 3: Performance degradation under Appearance Variations is mild. This is likely due to inherent redundancy in training data, which enhances robustness to facial variations. Models maintain high accuracy on accessory changes, contrasting with their vulnerability to age variations, indicating insufficient modeling of longitudinal facial changes.
@@
Loss Function Comparison. CosFace-IR and ArcFace-IR are tied as the top performers, achieving an accuracy of 97.17%. Their high clean accuracy suggests that these models effectively learn facial representations, enabling them to maintain capabilities when facing appearance changes.
Insight 4: While Corruption robustness shows weak correlation with clean accuracy, Variation robustness exhibits strong dependence. The decoupling of corruption resilience from clean performance likely stems from architectural limitations in handling structured noise. The high correlation for variations may stem from a domain adaptation effect, where models excel by enhancing feature disentanglement via margin-based losses like LMCL.